OpenStudy (anonymous):

The perimeter of a regular hexagon is seventy-two and each side is represented by two x. What is x? 6 9 twelve twenty-four

OpenStudy (anonymous):

6

OpenStudy (anonymous):

reg hexagon has 6 sides.. So each side is the same and totals up to 72. 72/6 = 12. Each side is 12. 12=2x x=6

OpenStudy (anonymous):

otays. thanks

OpenStudy (anonymous):

But howww?

OpenStudy (anonymous):

lol, what exactly dont you understand ?

OpenStudy (anonymous):

What part ?

OpenStudy (anonymous):

The perimeter of a regular hexagon is 72. A regular hexagon means each side is the same. A hexagon is 6-sided figure. So if there is 6 sides and they all are the same length & add up to 72. Perimeter(72)/6(sides)= each side is 12. & each side is represented by 2x. We know each side's length is 12. So: 2x=12 /2 /2 x=6

OpenStudy (anonymous):

Understand ?

OpenStudy (anonymous):

Yush

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok i know its one year ago but just saying the answer is 12 :)